OpenStudy (anonymous):

Why are most cells so small? In other words, what limits them from becoming larger? @abhisar

OpenStudy (aaronq):

Diffusion of materials across the membrane is the limiting factor. If we model the cell as a sphere, and compare the how the volume and the surface area of the sphere change with increasing radius we find that the volume increases much more rapidly. The large amount of "stuff" inside cannot be transported across the membrane at an acceptable after a certain radius size. Graph these and see what they look like. \(\sf V=\dfrac{4}{3}\pi r^3\) \(\sf SA=4\pi r^2\)